# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

### v2

@@ Appendix H Additional Details on the Relative Energy Computation and its Experimentation
[…]
which is the potential function $u({\mathbf{x}}_{t},t)$ of the generative process. Substituting this potential function (36) into the reverse SDE (2), we have the following generative process
$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (37)
For our experimental setting, where we use the DDPM [23] setting, we have the following variance preserving forward SDE
$\mathrm{d}{\mathbf{x}}_{t}=-\frac{1}{2}\beta(t){\mathbf{x}}_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}{\mathbf{w}}_{t}$ (38)
@@
$\mathrm{d}{\mathbf{x}}_{t}=\bigg{[}-\frac{1}{2}\beta(t){\mathbf{x}}_{t}-\beta(t)\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{]}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (39)
With Equation 37, we have
$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (40)
and by using Equation 35, the potential for our experimental setting equates to
[…]
@@
$\tilde{{\mathbf{x}}}(\alpha)=\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$ (42)
@@
$\tilde{u}(\alpha)=u(\tilde{{\mathbf{x}}}(\alpha))=\int^{\pi/2}_{0}\nabla u(\tilde{{\mathbf{x}}}(\alpha))\cdot{\mathbf{v}}\,\mathrm{d}\alpha$ (43)
where ${\mathbf{v}}=-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$. See Figure 22 for a visualization of this circular interpolation between two samples. For discrete times, we instead estimate the relative energy as
@@
given that $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ and the potential $u(\cdot)$ is obtained from the above Equation 41. Experimentation Details. In our experimentation of Section 7, we utilized a maximum of the top 2048 samples for each of three sample types (i.e., memorized, spurious, and generalized sets), and the training data set corresponding at each training data size $K$. Note, with the exception of the training data set, the samples in the three sets are sorted from least to greatest, based on their distance corresponding to their respective detection metric, e.g., Eq. (13), (14) or (15). We computed the relative energy of each set following Algorithm 2, which is strictly following the formulations detailed above. Note, we randomly selected and utilized an image from the entire training set as part of the relative energy calculation.
To define the average relative energy gap, let $\tilde{u}_{\text{target}}\in\mathbb{R}^{M}$ be a set of $M$ relative potentials computed from using either a memorized, spurious, or generalized set and following Equation 44. Please note that $M$ is capped at $2048$ samples. Since at some training data size $K$, certain sample type sets might not have less than $2048$ samples. For such cases, we use the entirety of that set to compute the relative energy of its samples. Meanwhile, let $\tilde{u}_{\text{data}}\in\mathbb{R}^{M^{\prime}}$ be the set of relative potentials computed from the training dataset at size $K$, where $M^{\prime}=2048$ for $K\geq 2048$. The average relative energy gap is defined as
@@
Using Equation 45, we computed the average relative energy gap and recorded our results in Figs. (8) and (23). The shaded regions reported in these figures represent the standard deviations computed from using $\tilde{u}_{\text{target}}$, and the average relative energy gap as the mean value. Since at some training data size $K$, there might be very few samples in one of the three sets, we opted not to include the relative energy gap of such sets, partly due to their large variance. Similar to the basin of attraction experiment, we determine whether a set is too small if the number of its samples is less than $0.1\%$ of its corresponding synthetic set. Finally, for our interpolation values $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ detailed in Algorithm 2 and used in Equation 44, we generate $20$ points starting from 0 to $\frac{\pi}{2}$ using linear spacing.
Moreover, since the relative energy computation requires a reference image, we studied the effects of having different reference images on the relative energy gap in Figs. (24), (25), (26), and (27). Overall, there is a slight difference in the values of the relative energy gap across the memorization-generalization transition for all datasets. However, the general trend is still maintained. Specifically, the gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential of spurious states quickly diverges as the later stages of generalization are initiated and consequently, where the gap between generalized and training data samples’ relative energy significantly decreases.
@@
H.1 Algorithm for Computing Relative Energy
1
2
3Inputs
4        Target pattern ${\mathbf{x}}_{1}\in\mathbb{R}^{N}$
5        Reference pattern ${\mathbf{x}}_{2}\in\mathbb{R}^{N}$
6        Score model $s_{\theta}$
7        Interpolation values $\bar{\alpha}$
8        Initial variance $\beta_{0}$
9Assumption: ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ are samples at $t=0$
10 Compute $\tilde{u}(\bar{\alpha}_{L})$
11        $\tilde{u}(\bar{\alpha}_{L})\leftarrow 0$
12
       $\bar{\alpha}\leftarrow\{0,\cdots,\frac{\pi}{2}\}$
@@ 11        $\tilde{u}(\bar{\alpha}_{L})\leftarrow 0$
13
14       $\Delta\alpha\leftarrow\bar{\alpha}_{1}-\bar{\alpha}_{0}$
15
16       for $\alpha\in\bar{\alpha}$ do
17               $\tilde{{\mathbf{x}}}\leftarrow\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$
18               $\tilde{{\mathbf{v}}}\leftarrow-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$
19               $\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\leftarrow-s_{\theta}(\tilde{{\mathbf{x}}},0)/\sqrt{\beta_{0}}$
               $\nabla_{\tilde{{\mathbf{x}}}}u(\tilde{{\mathbf{x}}})\cdot{\mathbf{v}}\,\Delta\alpha\leftarrow\big{(}-\frac{1}{2}\beta_{0}\tilde{{\mathbf{x}}}-\beta_{0}\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\big{)}\cdot{\mathbf{v}}\,\Delta\alpha$
                // See Equation 39
20
@@ 18               $\tilde{{\mathbf{v}}}\leftarrow-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$
                // See Equation 44
22
23
24Output
25        Relative energy $\tilde{u}(\bar{\alpha}_{L})$
@@
Algorithm 2 Compute the Discretized Relative Energy between a target ${\mathbf{x}}_{1}$ and a reference ${\mathbf{x}}_{2}$
H.2 Average Relative Energy Gap of Sample Types for All Datasets
@@
Figure 23: Average energy gap of each sample type and the training samples for the four datasets as the training data size $K$ grows, computed using Equation 44. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. With the exception of LSUN-CHURCH, the relative energy of the memorized samples matches that of the training data samples at $K=2$, but quickly diverges as $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
H.3 Average Relative Energy Gap using Different Reference Images for All Datasets
@@
Figure 24: Average energy gap of each sample type and the training samples of CIFAR10 as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
@@
Figure 25: Average energy gap of each sample type and the training samples of LSUN-CHURCH as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
@@
Figure 26: Average energy gap of each sample type and the training samples of FASHION-MNIST as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero (for three of the reference images) and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
@@
Figure 27: Average energy gap of each sample type and the training samples of MNIST as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero (for three of the reference images) and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.

### v3

@@ E.1  Overview
$u({\mathbf{x}}_{t},t)=g(t)^{2}\log p_{t}({\mathbf{x}}_{t})+\int^{0}_{{\mathbf{x}}_{t}}f({\mathbf{z}},t)d{\mathbf{z}},$ (39)
which is the potential function $u({\mathbf{x}}_{t},t)$ of the generative process. Substituting this potential function (39) into the reverse SDE, we have the following generative process:
$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}.$ (40)
For our experimental setting, where we used the DDPM [22] setting, we have the following variance preserving forward SDE:
$\mathrm{d}{\mathbf{x}}_{t}=-\frac{1}{2}\beta(t){\mathbf{x}}_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}{\mathbf{w}}_{t},$ (41)
@@
$\mathrm{d}{\mathbf{x}}_{t}=\bigg[-\frac{1}{2}\beta(t){\mathbf{x}}_{t}-\beta(t)\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg]\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}.$ (42)
With Eq. (40), we have
$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t},$ (43)
and by using Eq. (38), the potential for our experimental setting equates to
$\begin{split}u({\mathbf{x}}_{t},t)&=g(t)^{2}\log p_{t}({\mathbf{x}}_{t})+\int_{{\mathbf{x}}_{t}}^{0}f({\mathbf{z}},t)d{\mathbf{z}}\\ &=\beta(t)\log p_{t}({\mathbf{x}}_{t})+\frac{1}{2}\beta(t)\int_{{\mathbf{x}}_{t}}^{0}{\mathbf{z}}_{t}\mathrm{d}{\mathbf{z}}\\ &=\beta(t)\log p_{t}({\mathbf{x}}_{t})-\frac{1}{4}\beta(t){\mathbf{x}}_{t}^{2}.\end{split}$ (44)
@@
$\tilde{{\mathbf{x}}}(\alpha)=\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2},$ (45)
@@
$\tilde{u}(\alpha)=u(\tilde{{\mathbf{x}}}(\alpha))=\int^{\pi/2}_{0}\nabla u(\tilde{{\mathbf{x}}}(\alpha))\cdot{\mathbf{v}}\,\mathrm{d}\alpha,$ (46)
where ${\mathbf{v}}=-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$. See Fig. (25) for a visualization of this circular interpolation between two samples. For discrete times, we instead estimate the relative energy as
@@
given that $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ and the potential $u(\cdot)$ is obtained from the above Eq. (44).
E.2  Experimental Details
In our experimentation, we utilized a maximum of the top 512 samples for each of three sample types, memorized, spurious, and generalized, to compute the relative energy following Eq. (47), at each training data size $K$. For training data points, we instead utilized $\min(K,2048)$ number of samples. Note, with the exception of the training data set, the samples in the three sets are sorted from least to greatest, based on their distance corresponding to their respective detection metric (from the main text). We computed the relative energy of each set following Eq. (4), which is strictly following the formulations detailed above. Note, we randomly selected and utilized an image from the entire training set as part of the relative energy calculation.
To define the average relative energy gap, let $\tilde{u}_{\text{target}}\in\mathbb{R}^{M}$ be a set of $M$ relative potentials computed from using either a memorized, spurious, or generalized set and following Eq. (47). Please note that $M$ is capped at $512$ for the three sample types. Meanwhile, let $\tilde{u}_{\text{data}}\in\mathbb{R}^{M^{\prime}}$ be the set of relative potentials computed from the training dataset at size $K$ where $M^{\prime}=\min(K,2048)$. The average relative energy gap is defined as
@@ E.2  Experimental Details
Using Eq. (48), we computed the average relative energy gap and recorded our results in Fig. (26). The shaded regions reported in this figure represent the standard deviations computed from using $\tilde{u}_{\text{target}}$, and the average relative energy gap as the mean value. Since at some training data size $K$, there might be very few samples in one of the three sets, we opted not to include the relative energy gap of such sets, partly due to their large variance. Similar to the basin of attraction experiment, we determine whether a set is too small if the number of its samples is less than $0.1\%$ of its corresponding synthetic set. Finally, for our interpolation values $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ used for Eq. (47), we generate $20$ points starting from 0 to $\frac{\pi}{2}$ using linear spacing.
@@
E.3  Discussion
Since the relative energy computation requires a reference image, we also studied the effects of having different reference images on the relative energy gap in Figs. (27)-(30). Overall, the results for the relative energy gap across the memorization-generalization transition for all datasets are shown in Fig. (26). As the training data size $K$ increases, the gap for memorized samples is initially close to zero and quickly diverges. Spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase, where their potential is the lowest among the three sample types. However, the potential of spurious states quickly diverges as the later stages of generalization are initiated, and consequently, the gap between generalized and training data samples’ relative energies became the most narrowed among the three distinct sample types – further emphasizing the changing in the energy landscape as the DM undergoes its transition from memorization to generalization as $K$ increases.
Inputs
   Target pattern ${\mathbf{x}}_{1}\in\mathbb{R}^{N}$
  2.5pt Reference pattern ${\mathbf{x}}_{2}\in\mathbb{R}^{N}$
  2.5pt Score model $s_{\theta}$
  2.5pt Interpolation values $\bar{\alpha}$
  2.5pt Initial variance $\beta_{0}$
end
1exAssumption: ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ are samples at $t=0$
1exCompute $\tilde{u}(\bar{\alpha}_{L})$
  5pt $\bar{\alpha}\leftarrow\{0,\cdots,\frac{\pi}{2}\}$
@@ E.3  Discussion
  5pt
 $\Delta\alpha\leftarrow\bar{\alpha}_{1}-\bar{\alpha}_{0}$
  5pt
 $\tilde{u}(\bar{\alpha}_{L})\leftarrow 0$
  8pt
 for $\alpha\in\bar{\alpha}$ do
     5pt $\tilde{{\mathbf{x}}}\leftarrow\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$
     5pt $\tilde{{\mathbf{v}}}\leftarrow-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$
     5pt $\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\leftarrow-s_{\theta}(\tilde{{\mathbf{x}}},0)/\sqrt{\beta_{0}}$
     8pt $\nabla_{\tilde{{\mathbf{x}}}}u(\tilde{{\mathbf{x}}})\cdot{\mathbf{v}}\,\Delta\alpha\leftarrow\big(-\frac{1}{2}\beta_{0}\tilde{{\mathbf{x}}}-\beta_{0}\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\big)\cdot{\mathbf{v}}\,\Delta\alpha$
    // Eq. (42)
@@
    // Eq. (47)
   end for
end
Output
  5pt Relative energy $\tilde{u}(\bar{\alpha}_{L})$
end
@@
Algorithm 4 Compute the discretized relative energy between a target and a reference sample
E.4  Average Relative Energy Gap Results
@@
Figure 26: Average energy gap of each sample type and the training samples for the four datasets as the training data size $K$ grows, computed using Eq. (47). Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points. Shaded regions show standard deviation of the gap values. With the exception of LSUN-CHURCH, the relative energy of the memorized samples matches that of the training data samples at $K=2$, but quickly diverges as $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
E.5  Average Relative Energy Gap Results with Different References
@@
Figure 27: Average energy gap of each sample type and the training samples of CIFAR10 [37] as the training data size $K$ grows, computed using Eq. (47) and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
@@
Figure 28: Average energy gap of each sample type and the training samples of LSUN-CHURCH [75] as the training data size $K$ grows, computed using Eq. (47) and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
@@
Figure 29: Average energy gap of each sample type and the training samples of FASHION-MNIST [73] the training data size $K$ grows, computed using Eq. (47) and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero (for three of the reference images) and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
@@
Figure 30: Average energy gap of each sample type and the training samples of MNIST [17] as the training data size $K$ grows, computed using Eq. (47) and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero (for three of the reference images) and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
Appendix F Hardware Details
The training of the models detailed in Table (3) was done using NVIDIA Tesla V100 GPUs, where a single GPU was used to train each model. Each GPU has 32GB of memory and is linked with Power9 processors, clocking at 3.15 GHz maximum. Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively. Meanwhile, each LSUN-CHURCH model requires roughly 48, 56, and 103 hours of training with respect to the model’s initial latent dimension. In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model). Please note that these durations are with respect to the required training iterations for each model and their corresponding dataset detailed in Table (3). Also, note that we utilized the same hardware for our analyses.